# Modulation of Polarization and Metallicity in Janus Sliding Ferroelectrics

### v1

@@ Introduction
In recent years, the focus of the ferroelectric research community has shifted toward two-dimensional (2D) van der Waals (vdW) ferroelectrics for nanoscale device applications, owing to their atomic-scale thickness, clean surfaces, ease of integration, and high dielectric constants [1, 2]. Several promising candidates have been identified through first-principles simulations [3, 1]. Experimental demonstrations of ferroelectricity have included both in-plane polarization [4, 5, 6, 7, 8] and, in some materials, the more desirable out-of-plane polarization [9, 10, 11, 12, 13]. The coexistence of metallicity and switchable out-of-plane polarization remains exceptionally rare and has been observed primarily in semi-metallic 1T-WTe2 bilayer and multilayer systems [14]. To expand the functional landscape of low-dimensional ferroelectrics, new material designs are needed that offer out-of-plane polarization along with other useful features – such as a low bandgap for solar energy harvesting [15, 16], or metallic/semi-metallic behavior to explore quantum phenomena [17, 18, 19].
In 2017, Li et al. proposed an alternative mechanism for realizing vertical ferroelectric polarization in bilayers of certain 2D vdW materials by altering their stacking order through interlayer sliding [20]. This mechanism, known as sliding ferroelectricity [21], has since been experimentally confirmed in bilayers of hexagonal boron nitride (h-BN) [22, 23] and transition metal dichalcogenides (TMDs) [24, 25]. Owing to its interfacial nature, this stacking-engineered ferroelectric polarization can coexist with electrostatically doped in-plane conductivity [25, 26], opening avenues for engineering ferroelectric metals. Interestingly, a similar mechanism accounts for the coexistence of semi-metallicity and out-of-plane polarization in 1T-WTe2 [27]. Sliding ferroelectricity has also been demonstrated in MoS2 trilayers [25, 28, 29], where the presence of multiple polarization states offers prospects for designing multistate ferroelectric devices [28]. Due to the interfacial nature of sliding ferroelectrics, the polarization from both interfaces in trilayer systems contributes cumulatively, enhancing the overall polarization [25, 29]. More recently, first-principles calculations have predicted sliding ferroelectricity in Janus TMD bilayers [30, 31], further expanding the design landscape for 2D vdW ferroelectric materials by leveraging the unique properties of Janus structures [32].
@@
For the Janus systems, the layers are stacked such that the intrinsic electric field, $E$, of the Janus monolayers, will either point inward, i.e., towards the interface or will point outwards, i.e., towards the vacuum, as shown in Figure 1(b), resulting in E-in and E-out stacking for Janus bilayers. Other stacking orders, where the field $E$ points in the same direction in both the monolayers, will not result in sliding ferroelectricity in these Janus systems, as shown in a previous work [30], and hence are excluded in this study.
To evaluate the energetic stability of the Janus sliding ferroelectrics, we calculated the formation energy, $E_{F}$, for all the structures, as shown in Figure S1 [33]. The formation energy is calculated as the difference between the energy of the material system and the sum of the energies of the constituent atoms in their bulk phases. For the bilayers, this is given by,
@@
All the Janus bilayers show negative formation energies, confirming their energetic stability. We also compared the formation energies of the Janus bilayers to those of the experimentally synthesized systems (1T-WTe2 bilayer, Janus MoSSe monolayer, and AB-stacked WSe2 bilayer) to further confirm their stability, as shown in Figure S1 [33]. The crystal structure parameters and the formation energies for all the structures are provided in Table S1 [33].
@@
Figure 2: Polarization and interlayer charge density modulation in bilayer Janus sliding ferroelectrics.(a) Comparison of $P_{berry}$ (polarization via Berry phase method) for the Janus E-in (left) and E-out (right) bilayers. (b) $P_{berry}$ versus interlayer distance $d$ for XMoY bilayers. (c) Charge density difference profile for MoS2 bilayer. The yellow and blue colors represent the electron and hole accumulation, respectively. The dashed red lines represent the vertical location of the atoms. Black arrows denote the parameters $l_{h}$ (difference of hole accumulation peaks) and $l_{e}$ (electron accumulation peak) of the charge density profile. The variation of (d) $l_{h}$ and ($l_{e}$ parameters of the charge density profiles for the XMoY bilayers.
Next, we analyze and compare the polarization values of the Janus sliding ferroelectric bilayers, as shown in Figure 2(a). The polarization values are higher for E-in stacking and lower for E-out stacking of the Janus bilayers, suggesting polarization modulation due to the $E$ fields within the Janus monolayers. A similar polarization modulation was also observed in a previous work [31]. Along with polarization via the Berry phase method [34, 35], $P_{berry}$, we also calculated the interlayer potential difference, $\Delta\phi$, and polarization values, $P_{CD}$, from the electrostatic potential and the charge density profiles, respectively. Both $\Delta\phi$ and $P_{CD}$ exhibit modulation similar to $P_{berry}$ values for Janus bilayers, as shown in Table S2 [33].
The electrostatic potential and charge density profiles are obtained by calculating the differences between the plane-averaged potential and charge density, respectively, along the $z$-direction for the bilayer and the corresponding isolated monolayers. The difference between the electrostatic potential values obtained far (around 10Å) above the top and below the bottom surfaces of the bilayer provides $\Delta\phi$, similar to Ref. [25]. The methodology for calculating $P_{CD}$ value is provided in the Supplemental Material [33]. Using the electrostatic potential and charge density profiles, we can also show the polarization direction switching with a sliding motion, as shown in Figure S2 [33]. Note that the calculations of $\Delta\phi$ and $P_{CD}$ are computationally less demanding than that of $P_{berry}$ directly.
@@
To gain deeper insight into the modulation of interlayer distance and the resulting polarization, we further analyze the charge density profiles of the XMoY bilayers. The charge density profile of the MoS2 bilayer is shown in Figure 2(c), where two important features in the interlayer region of the plot are the difference between the hole accumulation peaks, $l_{h}$, and the height of the electron accumulation peak, $l_{e}$. The $l_{h}$ parameter can be used to quantify the charge redistribution between the layers because of the interlayer interaction, which is responsible for the magnitude of the out-of-plane polarization [31]. The $l_{e}$ parameter, as we have observed, correlates with the interlayer distance. The charge density profiles for all the XMoY bilayers are provided in Figure S3 [33] for comparison.
@@
The analysis of the charge density profiles for the structures with $d=d_{avg}$ explains how the modulation of the interlayer charge distribution via the E fields in the Janus sliding ferroelectrics results in the modulation of $d$. As seen in Figure 3(c), the E-in stacking reduces the electron accumulation between the bilayers by redistributing it towards the outer surface. As a result, the $l_{e}$ parameter decreases, as shown in Figure 3(d), which is expected to result in a reduction of the interlayer distance. Similarly, the E-out stacking results in an increased electron accumulation between the layers as shown in Figure S6[33] and depicted from the increase in $l_{e}$ in Figure 3(d), resulting in a larger interlayer distance on complete relaxation of the bilayers. Thus, the E fields within the Janus monolayers tune the electron accumulation within the interlayer region (or the $l_{e}$ parameter of the charge density profile), which modulates the interlayer distance of Janus sliding ferroelectrics.
Additionally, the higher values for $l_{h}$ in Figure S6 [33] for the XMoTe bilayers explain their higher polarization values for the case of $d=d_{avg}$. Thus, for the same $d$ values, the larger Y atoms result in stronger interlayer interactions, which explains the higher polarization values for XMoTe bilayers in Figure 3(a)-(b).
Finally, to validate our analysis, we repeated the polarization calculations – using $P_{CD}$ and $\Delta\phi$ – for the structures analyzed in Figure 3(b), varying the interlayer distance $d$ around $d_{avg}$, as presented in Figure S6 [33]. No effective polarization modulation was observed via the $E$ fields within Janus sliding ferroelectrics for fixed $d$.
@@
Following the polarization analysis, we examine the electronic band structures of the two Janus stacking configurations in Figure 4. The E-in stacking, shown in Figure 4(a), exhibits a lower bandgap than E-out stacking. This demonstrates that the bandgap can be tuned by altering the intrinsic electric field ($E$) configuration in Janus sliding ferroelectrics. Compared to the parent TMDs, the E-in stacked Janus structures exhibit reduced bandgaps, as shown in Table S2 [33]. This highlights a novel approach to achieving low-bandgap ferroelectrics through E-in stacking of Janus counterparts of TMD-based sliding ferroelectrics.
Next, we analyze the layer-resolved contributions to the valence and conduction bands to understand how the stacking sequence in Janus bilayers influences band structure modulation. We begin by comparing the layer-contribution-projected band structures of TeMoS (E-in stacking) and SMoTe (E-out stacking) bilayers, shown in Figure 4(b)-(c). At the valence band maximum (VBM), the TeMoS bilayer exhibits a more balanced contribution from the top and bottom layers than SMoTe. In contrast, at the conduction band minimum (CBM), the SMoTe bilayer shows a more equal layer contribution than TeMoS. Similar trends are observed across other XMoY bilayers, as illustrated in Figure S7 [33]. These results indicate that the $E$ fields within Janus monolayers modulates the layer-wise distribution of the VBM and CBM in Janus sliding ferroelectrics.
@@
VBM $\text{VBM ratio}=\frac{C_{top}^{\text{VBM}}}{C_{bottom}^{\text{VBM}}},$ (2)
CBM $\text{CBM ratio}=\frac{C_{bottom}^{\text{CBM}}}{C_{top}^{\text{CBM}}}.$ (3)
@@
To understand how these ratios are defined, we consider the direction of the interlayer electric field and its influence on the VBM and CBM contributions. The interlayer polarization points upward in the bilayers, as shown in Figure 1(a), resulting in a downward-pointing electric field, i.e., directed towards the bottom layer. This field increases the potential energy of electrons in the bottom layer, leading to a splitting of the VBM and CBM states. This effect is evident in the band structures of AB-stacked XMoY bilayers shown in Figure S7 [33], where the predominantly blue (bottom layer) bands appear at higher energies than the red (top layer) bands near the VBM and CBM. Consequently, the CBM and VBM show greater contributions from the top and bottom layers, respectively. A similar band alignment model near the Fermi level is proposed for bilayer sliding ferroelectrics in Ref. [26].
@@
We also investigated the effect of SOC on the band structures of the AB-stacked bilayers. Although SOC typically results in a slight reduction in the bandgap, it does not significantly alter the layer-wise contributions to the VBM and CBM – unless there is a shift in their positions. Notably, in MoTe2, WSe2, and SWSe bilayers, SOC causes the VBM to shift from the $\Gamma$ point to the K point, as shown in Figure S8 [33]. This shift leads to a pronounced decrease in the VBM ratio, due to the highly localized nature of the VBM at the K point, reflected by the dark-blue colored VBM.
Since the $E$ fields within Janus monolayers modulate the interlayer distance ($d$), we examine the band structures of E-in and E-out stacked Janus XMoY bilayers with $d$ fixed to $d_{avg}$, analogous to our earlier polarization analysis. For the constrained relaxation case, we find no significant difference in the bandgaps of the E-in and E-out configurations, as shown in Table S3 [33].
We further compare the VBM and CBM ratios for the XMoY bilayers with $d=d_{avg}$ in Table S3[33]. The E-in stacked Janus bilayers exhibit higher VBM ratios than their E-out counterparts. However, due to a shift in the CBM position for these Janus bilayers, a direct comparison of the CBM ratios is less straightforward.
Overall, fixing $d$ eliminates the bandgap variation due to the $E$ fields. However, variations in the layer-wise contributions to the VBM and CBM remain observable. This indicates that, as with polarization, the bandgap modulation in Janus bilayers due to the $E$ fields primarily arises from the tuning of $d$. Nevertheless, the direction of the $E$ fields in the Janus bilayers, i.e., E-in or E-out stacking, continues to play a crucial role in determining the layer-resolved contributions to the electronic bands.
@@
As shown in Refs. [25, 26], the extent of polarization reduction caused by extrinsic charge doping depends on the relative contributions of each layer to the valence and conduction bands. When both layers contribute equally, the polarization reduction or the depolarization effect from electron or hole doping is minimized. Since the intrinsic electric fields ($E$) within the Janus monolayers modulate the VBM and CBM contributions in the Janus bilayers, they can also modulate the doping-induced depolarization. This, in turn, allows modulation of the coexistence of ferroelectric polarization and dopant-induced metallicity.
The effect of charge dopants on the polarization can be understood by comparing the reduction in the magnitude of polarization with charge doping relative to the undoped bilayer system. This phenomenon of polarization reduction is referred to as ’depolarization’ [25]. Note that instead of polarization, we calculated the variation in the interlayer potential difference ($\Delta\phi$) with charge doping, as shown in Figure S9 and S10 [33], as polarization for metallic systems is not well-defined, following the approach used in previous studies [25, 26]. Using this definition, the percentage of depolarization due to hole doping (Depolhole) and electron doping (Depolelectron), are calculated as
@@
Although one can not define polarization for metallic systems as per the Berry phase method [34, 35], we can get an estimate of the polarization from $P_{CD}$ values, as shown in Figure S9 [33]. $P_{CD}$ shows a similar doping-induced modulation as $\Delta\phi$.
@@
Since the doped holes and electrons populate the valence and conduction bands, respectively, their distribution in the reciprocal space is influenced by the layer-wise contribution of the VBM and CBM. This, in turn, affects the layer-wise distribution of charge carriers and the resulting charge imbalance between layers, which drives polarization, as shown in Figure S2(b) [33]. Note that a higher VBM or CBM ratio results in a more uniformly layer-wise distributed charge density at the VBM and CBM, respectively, as can be seen from the partial charge density plots at the VBM and CBM in Figure 4(b)-(c) and also in Figure S7 [33]. This uniformity reduces the impact of doped carriers on charge redistribution and thus on polarization. A more detailed analysis of the VBM and CBM layer-wise contribution and its effect on the doping-induced depolarization can be found in Ref. [26].
We also examined the effect of SOC on the variation of $\Delta\phi$ with charge doping, as shown in Figures S9 and S10 [33]. Notable differences between SOC and non-SOC results were observed only for hole doping in MoTe2, WSe2, and SWSe bilayers, which can be attributed to the SOC-induced shift in the VBM, as shown in Figure S8 [33].
@@
We have identified the tuning of the interlayer distance ($d$) as the mechanism behind the polarization and bandgap modulation in Janus sliding ferroelectrics. In this section, we explicitly calculate the effect of the variation of $d$ on the polarization and band structure of the bilayer sliding ferroelectrics.

### v2

@@ Introduction
In recent years, the focus of the ferroelectric research community has shifted toward two-dimensional (2D) van der Waals (vdW) ferroelectrics for nanoscale device applications, owing to their atomic-scale thickness, clean surfaces, ease of integration, and high dielectric constants [Xiao2019, MenghaoWu2021]. Several promising candidates have been identified through first-principles simulations [Kruse2023, Xiao2019]. Experimental demonstrations of ferroelectricity have included both in-plane polarization [Chang2016, Chang2020, Higashitarumizu2020, Ghosh2019, Gou2023] and, in some materials, the more desirable out-of-plane polarization [Liu2016, Belianinov2015, Cui2018, Zhou2017, Yuan2019]. The coexistence of metallicity and switchable out-of-plane polarization remains exceptionally rare and has been observed primarily in semi-metallic 1T-WTe2 bilayer and multilayer systems [Fei2018]. To expand the functional landscape of low-dimensional ferroelectrics, new material designs are needed that offer out-of-plane polarization along with other useful features – such as a low bandgap for solar energy harvesting [Zhang2018, Wu2022], or metallic/semi-metallic behavior to explore quantum phenomena [Bhowal2023, Zhou2020, Young2023].
In 2017, Li et al. proposed an alternative mechanism for realizing vertical ferroelectric polarization in bilayers of certain 2D vdW materials by altering their stacking order through interlayer sliding [Li2017]. This mechanism, known as sliding ferroelectricity [Menghao2021], has since been experimentally confirmed in bilayers of hexagonal boron nitride (h-BN) [Yasuda2021, Stern2021] and transition metal dichalcogenides (TMDs) [Wang2022, Deb2022]. Owing to its interfacial nature, this stacking-engineered ferroelectric polarization can coexist with electrostatically doped in-plane conductivity [Deb2022, Wang2024], opening avenues for engineering ferroelectric metals. Interestingly, a similar mechanism accounts for the coexistence of semi-metallicity and out-of-plane polarization in 1T-WTe2 [Yang2018]. Sliding ferroelectricity has also been demonstrated in MoS2 trilayers [Deb2022, Meng2022, Cao2024], where the presence of multiple polarization states offers prospects for designing multistate ferroelectric devices [Meng2022]. Due to the interfacial nature of sliding ferroelectrics, the polarization from both interfaces in trilayer systems contributes cumulatively, enhancing the overall polarization [Deb2022, Cao2024]. More recently, first-principles calculations have predicted sliding ferroelectricity in Janus TMD bilayers [Lin2024, Zhang2024], further expanding the design landscape for 2D vdW ferroelectric materials by leveraging the unique properties of Janus structures [Zhang2020].
@@
Computational Methods
Our first-principles calculations were carried out within the density functional theory (DFT) formalism, as implemented in the quantum espresso code [Giannozzi2017, Giannozzi2009]. We used ultrasoft pseudopotentials [Vanderbilt1990] with a kinetic energy cut-off of 90 Ry for the wavefunctions and 720 Ry for charge density, with a $k$-mesh of $14\times 14\times 1$. We employed the Perdew-Burke-Ernzerhof (PBE) formalism to describe the exchange-correlation functional [Perdew1996], along with the Grimme-D3 dispersion correction with BJ damping [Grimme2010]. We used a vacuum of around 30Åto avoid interaction between adjacent images due to the periodic boundary conditions. The electronic self-consistent cycle used the convergence criterion of $1\times 10^{-8}$ eV. We relaxed the structures until the residual forces on the atoms were less than 0.001 eV/Å. We calculated the electric polarization values using the Berry phase method [King1993, Resta1994], unless specified otherwise. The electrostatic potential profiles were obtained using the ionic and Hartree contributions from the charge density [Wang2022, Ferreira2021]. We also used a dipole moment correction along the out-of-plane direction for the electrostatic potential profiles.
We implemented charge doping calculations using the fractional nuclear charge pseudo atom approach [Sinai2013]. Following a similar approach as Ref. [Deb2022], we carried out the doping calculations by charging the metal nuclei. The charging of the metal nuclei was done by adding an equal amount of fractional charge to the metal nuclei of both layers for all the bilayer systems. The maximum doping limit for electron and hole dopants we used in this study is $\sim$1013 cm-2, which is the highest reported experimentally accessible charge density in such systems [Deb2022]. At each doping stage, we relaxed the bilayers keeping the lattice constants fixed. We used Fermi-Dirac smearing for the band structure and density of states (DOS) calculations and for the self-consistent cycle during the charge doping to enhance the convergence. We employed spin-orbit interactions for band structure, DOS, and charge doping calculations, as well as other cases as explicitly mentioned in the results section. We carried out phonon calculations within the finite displacement approach using a $2\times 2\times 1$ supercell, as implemented using a combination of quantum espresso and phonopy codes [phonopy-phono3py-JPCM, phonopy-phono3py-JPSJ].
@@
For the Janus systems, the layers are stacked such that the intrinsic electric field, $E$, of the Janus monolayers, will either point inward, i.e., towards the interface or will point outwards, i.e., towards the vacuum, as shown in Figure 1(b), resulting in E-in and E-out stacking for Janus bilayers. Other stacking orders, where the field $E$ points in the same direction in both the monolayers, will not result in sliding ferroelectricity in these Janus systems, as shown in a previous work [Lin2024], and hence are excluded in this study.
To evaluate the energetic stability of the Janus sliding ferroelectrics, we calculated the formation energy, $E_{F}$, for all the structures, as shown in Figure S1 [SuppMat]. The formation energy is calculated as the difference between the energy of the material system and the sum of the energies of the constituent atoms in their bulk phases. For the bilayers, this is given by,
@@
All the Janus bilayers show negative formation energies, confirming their energetic stability. We also compared the formation energies of the Janus bilayers to those of the experimentally synthesized systems (1T-WTe2 bilayer, Janus MoSSe monolayer, and AB-stacked WSe2 bilayer) to further confirm their stability, as shown in Figure S1 [SuppMat]. The crystal structure parameters and the formation energies for all the structures are provided in Table S1 [SuppMat].
@@
[…]
Next, we analyze and compare the polarization values of the Janus sliding ferroelectric bilayers, as shown in Figure 2(a). The polarization values are higher for E-in stacking and lower for E-out stacking of the Janus bilayers, suggesting polarization modulation due to the $E$ fields within the Janus monolayers. In general, the E-in Janus counterparts of TMD bilayers exhibit enhanced polarization, whereas the E-out stacked Janus bilayers show reduced polarization compared to their parent TMD bilayers, as summarized in Table S2 [SuppMat]. A similar polarization modulation was also observed in a previous work [Zhang2024]. Along with polarization via the Berry phase method [King1993, Resta1994], $P_{berry}$, we also calculated the interlayer potential difference, $\Delta\phi$, and polarization values, $P_{CD}$, from the electrostatic potential and the charge density profiles, respectively. Both $\Delta\phi$ and $P_{CD}$ exhibit modulation similar to $P_{berry}$ values for Janus bilayers, as shown in Table S2 [SuppMat].
The electrostatic potential and charge density profiles are obtained by calculating the differences between the plane-averaged potential and charge density, respectively, along the $z$-direction for the bilayer and the corresponding isolated monolayers. The difference between the electrostatic potential values obtained far (around 10Å) above the top and below the bottom surfaces of the bilayer provides $\Delta\phi$, similar to Ref. [Deb2022]. The methodology for calculating $P_{CD}$ value is provided in the supplementary material [SuppMat]. Using the electrostatic potential and charge density profiles, we can also show the polarization direction switching with a sliding motion, as shown in Figure S2 [SuppMat]. Note that the calculations of $\Delta\phi$ and $P_{CD}$ are computationally less demanding than that of $P_{berry}$ directly.
@@
To gain deeper insight into the modulation of interlayer distance and the resulting polarization, we further analyze the charge density profiles of the XMoY bilayers. The charge density profile of the MoS2 bilayer is shown in Figure 2(c), where two important features in the interlayer region of the plot are the difference between the hole accumulation peaks, $l_{h}$, and the height of the electron accumulation peak, $l_{e}$. The $l_{h}$ parameter can be used to quantify the charge redistribution between the layers because of the interlayer interaction, which is responsible for the magnitude of the out-of-plane polarization [Zhang2024]. The $l_{e}$ parameter, as we have observed, correlates with the interlayer distance. The charge density profiles for all the XMoY bilayers are provided in Figure S3 [SuppMat] for comparison.
@@
To assess the role of relativistic effects on out-of-plane polarization, we compared the charge density and electrostatic potential profiles of XMoY bilayers with and without spin–orbit coupling (SOC), as shown in Figure S3 and Figure S4 [SuppMat], respectively. The profiles exhibit no significant differences, indicating minimal impact of SOC on ferroelectric polarization. A quantitative comparison of $\Delta\phi$ and $P_{CD}$ values, presented in Figure S5 [SuppMat], further confirms that SOC introduces only negligible changes, and thus does not substantially affect the switchable out-of-plane polarization in these systems.
@@
The analysis of the charge density profiles for the structures with $d=d_{avg}$ explains how the modulation of the interlayer charge distribution via the E fields in the Janus sliding ferroelectrics results in the modulation of $d$. As seen in Figure 3(c), the E-in stacking reduces the electron accumulation between the bilayers by redistributing it towards the outer surface. As a result, the $l_{e}$ parameter decreases, as shown in Figure 3(d), which is expected to result in a reduction of the interlayer distance. Similarly, the E-out stacking results in an increased electron accumulation between the layers as shown in Figure S6[SuppMat] and depicted from the increase in $l_{e}$ in Figure 3(d), resulting in a larger interlayer distance on complete relaxation of the bilayers. Thus, the E fields within the Janus monolayers tune the electron accumulation within the interlayer region (or the $l_{e}$ parameter of the charge density profile), which modulates the interlayer distance of Janus sliding ferroelectrics.
Additionally, the higher values for $l_{h}$ in Figure S6 [SuppMat] for the XMoTe bilayers explain their higher polarization values for the case of $d=d_{avg}$. Thus, for the same $d$ values, the larger Y atoms result in stronger interlayer interactions, which explains the higher polarization values for XMoTe bilayers in Figure 3(a)-(b).
Finally, to validate our analysis, we repeated the polarization calculations – using $P_{CD}$ and $\Delta\phi$ – for the structures analyzed in Figure 3(b), varying the interlayer distance $d$ around $d_{avg}$, as presented in Figure S6 [SuppMat]. No effective polarization modulation was observed via the $E$ fields within Janus sliding ferroelectrics for fixed $d$.
@@
Following the polarization analysis, we examine the electronic band structures of the two Janus stacking configurations in Figure 4. The E-in stacking, shown in Figure 4(a), exhibits a lower bandgap than E-out stacking. This demonstrates that the bandgap can be tuned by altering the intrinsic electric field ($E$) configuration in Janus sliding ferroelectrics. Compared to the parent TMDs, the E-in stacked Janus structures exhibit reduced bandgaps, as shown in Table S2 [SuppMat]. This highlights a novel approach to achieving low-bandgap ferroelectrics through E-in stacking of Janus counterparts of TMD-based sliding ferroelectrics.
Next, we analyze the layer-resolved contributions to the valence and conduction bands to understand how the stacking sequence in Janus bilayers influences band structure modulation. We begin by comparing the layer-contribution-projected band structures of TeMoS (E-in stacking) and SMoTe (E-out stacking) bilayers, shown in Figure 4(b)-(c). At the valence band maximum (VBM), the TeMoS bilayer exhibits a more balanced contribution from the top and bottom layers than SMoTe. In contrast, at the conduction band minimum (CBM), the SMoTe bilayer shows a more equal layer contribution than TeMoS. Similar trends are observed across other XMoY bilayers, as illustrated in Figure S7 [SuppMat]. These results indicate that the $E$ fields within Janus monolayers modulates the layer-wise distribution of the VBM and CBM in Janus sliding ferroelectrics.
@@
$\text{VBM ratio}=\frac{C_{top}^{\text{VBM}}}{C_{bottom}^{\text{VBM}}},$ (2)
$\text{CBM ratio}=\frac{C_{bottom}^{\text{CBM}}}{C_{top}^{\text{CBM}}}.$ (3)
@@
To understand how these ratios are defined, we consider the direction of the interlayer electric field and its influence on the VBM and CBM contributions. The interlayer polarization points upward in the bilayers, as shown in Figure 1(a), resulting in a downward-pointing electric field, i.e., directed towards the bottom layer. This field increases the potential energy of electrons in the bottom layer, leading to a splitting of the VBM and CBM states. This effect is evident in the band structures of AB-stacked XMoY bilayers shown in Figure S7 [SuppMat], where the predominantly blue (bottom layer) bands appear at higher energies than the red (top layer) bands near the VBM and CBM. Consequently, the CBM and VBM show greater contributions from the top and bottom layers, respectively. A similar band alignment model near the Fermi level is proposed for bilayer sliding ferroelectrics in Ref. [Wang2024].
@@
We also investigated the effect of SOC on the band structures of the AB-stacked bilayers. Although SOC typically results in a slight reduction in the bandgap, it does not significantly alter the layer-wise contributions to the VBM and CBM – unless there is a shift in their positions. Notably, in MoTe2, WSe2, and SWSe bilayers, SOC causes the VBM to shift from the $\Gamma$ point to the K point, as shown in Figure S8 [SuppMat]. This shift leads to a pronounced decrease in the VBM ratio, due to the highly localized nature of the VBM at the K point, reflected by the dark-blue colored VBM.
Since the $E$ fields within Janus monolayers modulate the interlayer distance ($d$), we examine the band structures of E-in and E-out stacked Janus XMoY bilayers with $d$ fixed to $d_{avg}$, analogous to our earlier polarization analysis. For the constrained relaxation case, we find no significant difference in the bandgaps of the E-in and E-out configurations, as shown in Table S3 [SuppMat].
We further compare the VBM and CBM ratios for the XMoY bilayers with $d=d_{avg}$ in Table S3[SuppMat]. The E-in stacked Janus bilayers exhibit higher VBM ratios than their E-out counterparts. However, due to a shift in the CBM position for these Janus bilayers, a direct comparison of the CBM ratios is less straightforward.
Overall, fixing $d$ eliminates the bandgap variation due to the $E$ fields. However, variations in the layer-wise contributions to the VBM and CBM remain observable. This indicates that, as with polarization, the bandgap modulation in Janus bilayers due to the $E$ fields primarily arises from the tuning of $d$. Nevertheless, the direction of the $E$ fields in the Janus bilayers, i.e., E-in or E-out stacking, continues to play a crucial role in determining the layer-resolved contributions to the electronic bands. Thus, E-field–based stacking of Janus sliding ferroelectrics offers an additional strategy for band structure modulation in 2D materials, complementing the more commonly employed approach of constructing versatile heterostructures [Nguyen2021, Truong2025, Nguyen2025].
@@
As shown in Refs. [Deb2022, Wang2024], the extent of polarization reduction caused by extrinsic charge doping depends on the relative contributions of each layer to the valence and conduction bands. When both layers contribute equally, the polarization reduction or the depolarization effect from electron or hole doping is minimized. Since the intrinsic electric fields ($E$) within the Janus monolayers modulate the VBM and CBM contributions in the Janus bilayers, they can also modulate the doping-induced depolarization. This, in turn, allows modulation of the coexistence of ferroelectric polarization and dopant-induced metallicity.
The effect of charge dopants on the polarization can be understood by comparing the reduction in the magnitude of polarization with charge doping relative to the undoped bilayer system. This phenomenon of polarization reduction is referred to as ’depolarization’ [Deb2022]. Note that instead of polarization, we calculated the variation in the interlayer potential difference ($\Delta\phi$) with charge doping, as shown in Figure S9 and S10 [SuppMat], as polarization for metallic systems is not well-defined, following the approach used in previous studies [Deb2022, Wang2024]. Using this definition, the percentage of depolarization due to hole doping (Depolhole) and electron doping (Depolelectron), are calculated as
@@
Although one can not define polarization for metallic systems as per the Berry phase method [King1993, Resta1994], we can get an estimate of the polarization from $P_{CD}$ values, as shown in Figure S9 [SuppMat]. $P_{CD}$ shows a similar doping-induced modulation as $\Delta\phi$.
@@
To clarify the conditions under which metallicity and polarization coexist in doped bilayers, we examined the variation of the Fermi energy with doping and compared the density of states (DOS) of undoped and doped TeMoS and SMoTe bilayers, as shown in Figures S11 and S12 [SuppMat]. Electron (hole) doping raises (lowers) the Fermi energy, eventually shifting it across the CBM (VBM) of the undoped bilayer. For the highest doping levels, the Fermi energy falls within regions of non-zero DOS in the undoped case, and the corresponding DOS plots for doped bilayers confirm finite DOS values at the Fermi level, indicating sufficient free charge carriers for metallic behavior. Similar results are observed across other bilayers, where doping near $10^{13}$ cm-2 consistently yields non-zero DOS at the Fermi level. Hence, bilayers doped at this level are expected to exhibit appreciable in-plane conductivity alongside switchable out-of-plane polarization, thereby fulfilling the criteria of a ferroelectric metal.
Since the doped holes and electrons populate the valence and conduction bands, respectively, their distribution in the reciprocal space is influenced by the layer-wise contribution of the VBM and CBM. This, in turn, affects the layer-wise distribution of charge carriers and the resulting charge imbalance between layers, which drives polarization, as shown in Figure S2(b) [SuppMat]. Note that a higher VBM or CBM ratio results in a more uniformly layer-wise distributed charge density at the VBM and CBM, respectively, as can be seen from the partial charge density plots at the VBM and CBM in Figure 4(b)-(c) and also in Figure S7 [SuppMat]. This uniformity reduces the impact of doped carriers on charge redistribution and thus on polarization.
In the E-in stacked Janus TeMoS bilayer, the CBM charge density is localized on the top layer, whereas the VBM charge density is more evenly distributed across both layers [Figure 4(b)]. As a result, doped electrons preferentially occupy the top layer, leading to stronger electron-induced depolarization, while doped holes are redistributed more uniformly, producing weaker hole-doping-induced depolarization. Reversing the intrinsic electric field, as in the E-out stacked Janus SMoTe bilayer, inverts the nature of charge density profiles of both the VBM and CBM. In this case, the VBM charge density localizes on the bottom layer [Figure 4(c)], giving rise to stronger hole-doping-induced depolarization. A more detailed analysis of the VBM and CBM layer-wise contribution and its effect on the doping-induced depolarization can be found in Ref. [Wang2024].
We also examined the effect of SOC on the variation of $\Delta\phi$ with charge doping, as shown in Figures S9 and S10 [SuppMat]. Notable differences between SOC and non-SOC results were observed only for hole doping in MoTe2, WSe2, and SWSe bilayers, which can be attributed to the SOC-induced shift in the VBM, as shown in Figure S8 [SuppMat].
@@
We have identified the tuning of the interlayer distance ($d$) as the mechanism behind the polarization and bandgap modulation in Janus sliding ferroelectrics. A widely used approach to tune the interlayer distance in bilayer and heterostructured TMDs is the application of out-of-plane hydrostatic pressure using a diamond anvil cell [Qiao2022, Li2023]. In addition, in-plane strain engineering has also been shown to effectively modulate the interlayer coupling [Pak2017, Yan2020]. In this section, we explicitly calculate the effect of the variation of $d$ on the polarization and band structure of the bilayer sliding ferroelectrics.